undeadknight26 (undeadknight26):

The graph of f(x) = (0.5)x is replaced by the graph of g(x) = (0.5)x - k. If g(x) is obtained by shifting f(x) down by 7 units, the value of k is ________.

undeadknight26 (undeadknight26):

@ash2326 @skullpatrol please help

OpenStudy (solomonzelman):

g(x) = (0.5)x - k is shifted k units down from f(x) = (0.5)x

undeadknight26 (undeadknight26):

oh my bad its: g(x) = (.5)^x and f(x) = (.5)^x.

undeadknight26 (undeadknight26):

WOuld you just plug in 7 for x?

OpenStudy (solomonzelman):

If g(x) is obtained by shifting f(x) down by 7 units meaning that g(x)=(0.5)~~~~~7 units to the bottom from f(x)-k? I can't explain but k=-7 it's g(x)=f(x)-(-7) so the k=7. The question is corrected can you rephrase it?

undeadknight26 (undeadknight26):

so the answer is 7?

OpenStudy (solomonzelman):

It's -7 f(x)-k=f(g)-7 the original function of f(x) is obtained by shifting f(g) 7 units down, or by -7 so x=7 YES!

undeadknight26 (undeadknight26):

oh ok thank you solo!

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

@undeadknight26 do you remember if you got this right?

undeadknight26 (undeadknight26):

if solo showed me it...its right :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks